Condensed balance sheet and income statement data for Jergan Corporation are presented here.

Jergan Corporation
Balance Sheets
December 31

2020

2019

2018

Cash $ 30,100 $ 17,500 $ 18,100
Accounts receivable (net) 49,000 45,400 47,100
Other current assets 89,500 95,400 64,400
Investments 54,400 70,300 45,700
Plant and equipment (net) 499,200 370,100 357,100
$722,200 $598,700 $532,400
Current liabilities $84,400 $80,300 $69,100
Long-term debt 144,500 86,000 50,100
Common stock, $10 par 321,000 317,000 297,000
Retained earnings 172,300 115,400 116,200
$722,200 $598,700 $532,400

Jergan Corporation
Income Statement
For the Years Ended December 31

2020

2019

Sales revenue $737,500 $607,000
Less: Sales returns and allowances 40,500 30,600
Net sales 697,000 576,400
Cost of goods sold 426,400 371,000
Gross profit 270,600 205,400
Operating expenses (including income taxes) 186,960 153,524
Net income $ 83,640 $ 51,876


Additional information:

1. The market price of Jergan’s common stock was $7.00, $7.50, and $8.50 for 2018, 2019, and 2020, respectively.
2. You must compute dividends paid. All dividends were paid in cash.


(a)

Compute the following ratios for 2019 and 2020. (Round Asset turnover and Earnings per share to 2 decimal places, e.g. 1.65. Round payout ratio and debt to assets ratio to 0 decimal places, e.g. 18%. Round all other answers to 1 decimal place, e.g. 6.8 or 6.8%.)

2019

2020

(1) Profit margin % %
(2) Gross profit rate % %
(3) Asset turnover times times
(4) Earnings per share $ $
(5) Price-earnings ratio times times
(6) Payout ratio % %
(7) Debt to assets ratio % %
Answer #1
2019 Working-2019 2020 Working-2020
1 Profit margin 8.5% =51876/607000 11.3% =83640/737500
2 Gross profit rate 33.8% =205400/607000 36.7% =270600/737500
3 Asset turnover         1.07 =607000/((598700+532400)/2) 1.70 =737500/((499200+370100)/2)
4 Earnings per share         1.64 =51876/(317000/10) 2.61 =83640/(321000/10)
5 Price-earnings ratio          4.6 =7.5/1.64     3.3 =8.5/2.61
6 Payout ratio 101% =((116200+51786-115400)/(317000/10))/1.64 32% =((115400+83640-172300)/(321000/10))/2.61
7 Debt to assets ratio 28% =(80300+86000)/598700 32% =(84400+144500)/722200